OpenStudy (anonymous):

Don randomly draws two cards from a standard deck of 52 cards. He does not replace the first card. What is the probability that both cards are aces?

OpenStudy (kropot72):

The probability of an ace on the first draw is 4/52. If an ace is drawn on the first draw, the are 3 aces remaining an 51 cards in the pack. So the probability of an ace on the second draw is 3/51. The required probability is therefore given by: \[\frac{4}{52}\times\frac{3}{51}=you\ can\ calculate\]

OpenStudy (anonymous):

thx a bunch